OpenStudy (anonymous):

Convert 2rad3 +2i into its polar form

OpenStudy (anonymous):

2rad3?

OpenStudy (anonymous):

\[2\sqrt 3 +2i\] is this your problem?

OpenStudy (anonymous):

Let z=2√3+2i comparing it with z=a+bi we find a= 2√3, b= 2 Now \[r=\sqrt{a^2+b^2} = \sqrt{(2\sqrt3)^2+2^2}= \sqrt{12+6}= \sqrt{16} = 4\] \[\theta = \tan^{-1} (\frac{b}{a}) =\tan^{-1} (\frac{2}{2\sqrt3})=\tan^{-1} (\frac{1}{\sqrt3}) =30^0\] Polar form of a complex number is given as \[z=r(\cos \theta + i \sin \theta) \] i.e. \[z=4(\cos30^0 + i \sin 30^0) \] Which is the required polar form of the given complex number z=2√3+2i

OpenStudy (anonymous):

Sorry, in the fourth line write 4 instead of 6